Find the interval of convergence of the power series 0 - က၊ 8 2. [ 9 - 73n + 5) ၏ 0 ၀ ၊ () { 82 } က 0 ကံ 0 8 8 05. [-8,
Answer #1

* Givery series Here an = (3nts) Radius of convergence = linken la k- (ཀ = licerents) - ling om Series egs (y 1x28h 6 - 8/3 23nts = 8 (2nts divergen 7- in of n=813, then & ( 3nts ) 2 by nth term test so, the series diverges at end points . -8,8h Henc